# FuzzyLight: A Robust Two-Stage Fuzzy Approach for Traffic Signal Control Works in Real Cities

### v1

@@ 2.1. Traffic Signal Control
A traffic signal phase comprises a set of permitted traffic movements. As depicted in Figure 2(c), we consider $12$ traffic movements, and $p_{i}$ represnts the $i-th$ TSP . In this configuration, we use four distinct TSP phases.
@@ 3.1.1. Fuzzification: membership function with compressed sensing
Initially, to meet the requirements of compressed sensing, a negotiation between the sensor and the traffic signal controller establishes a random matrix $A\in\mathbb{R}^{z\times n}$. Following this, we construct the sensor data into a sparse matrix $X\in\mathbb{R}^{n\times m}$ to fulfill the sparsity condition essential for effective compressed sensing. It represents the number of waiting vehicles in all lanes, where $n$ represents the number of entry lanes at the intersection $\mathcal{L}^{in}$, and $m=\frac{ER}{slice}$. In this matrix, $slice=v_{l}+safe_{l}$, where $v_{l}$ is the vehicle queue length, enhancing the sparsity, and $safe_{l}$ is the safety distance between vehicles. The $ER$ denotes the effective observation range used to calibrate the sensor deployment, optimizing cost and performance.
@@ 3.1.2. Defuzzification: denoising function and fuzzy rules for TSP
where $round$ signifies rounding up to the nearest integer. Based on $X_{defuzi}$, the number of waiting vehicles for each phase $WV_{p_{i}}$ is calculated, and fuzzy rules (detailed in Algorithm 1 lines 7-9) are used to determine the phase $p_{dynamic}$ corresponding to the largest number of waiting vehicles.
@@ 3.2.1. Fuzzification: Neural network fuzzy membership function (NNFM)
Lane Embedding Layer: First, we extract the number of vehicles of different segments ($seg=k*slice$) of each traffic movement by recovered vehicle matrix $\hat{X}$. Where $k$ defines the length of the segment within each lane. To allow more capacity in feature extraction, each segment of the lane is embedded from one-dimensional into a higher dimensional latent space using a layer of multi-layer perception (MLP):
@@
(12) $t_{duration}=clip\left(h_{fuzzi}\ast refer\ duration+\epsilon,low,high\right)\$
Where the $OU$ noise (Lillicrap et al., 2015) $\epsilon$ is used for enhancing phase duration exploration. The $refer$ $duration$ is calculated by maximum lane length $MaxLaneLength$ divided by maximum lane speed $MaxLaneSpeed$: $refer\ duration=\frac{MaxLaneLength}{MaxLaneSpeed}$.
We clip the phase duration to a maximum and minimum value to guarantee safety. To ensure pedestrian safety, set the pedestrian crossing time as the minimum duration of green lights $low$.
@@
5:     for Each episode in $epochs$ do
@@ 5. Discussion
Scale (Max_scale) JiNan HangZhou
@@
Scale (Max_scale) JiNan HangZhou
@@ 2. How the FuzzyLight affects the throughput rate?
where $t_{duration}=\frac{dis}{v}$, $dis$ represents the distance of the last vehicle from intersections and $v$ denotes the average speed of the vehicle. We Then the throughput rate of a phase pair can be expressed as
@@
if $x(l_{i})=x(l_{j})$ then $Throughput=2$, therefore $21<=Throughput<=$ 2.

### v2

@@ 2.1. Traffic Signal Control
A traffic signal phase comprises a set of permitted traffic movements. As depicted in Figure 2(c), we consider $12$ traffic movements, and $p_{i}$ represnts the ithi-th TSP . In this configuration, we use four distinct TSP phases.
@@ 3.1.1. Fuzzification: membership function with compressed sensing
Initially, to meet the requirements of compressed sensing, a negotiation between the sensor and the traffic signal controller establishes a random matrix $A\in\mathbb{R}^{z\times n}$. Following this, we construct the sensor data into a sparse matrix $X\in\mathbb{R}^{n\times m}$ to fulfill the sparsity condition essential for effective compressed sensing. It represents the number of waiting vehicles in all lanes, where $n$ represents the number of entry lanes at the intersection $\mathcal{L}^{in}$, and $m=\frac{ER}{slice}$. In this matrix, $slice=v_{l}+safe_{l}$, where $v_{l}$ is the vehicle length, enhancing the sparsity, and $safe_{l}$ is the safety distance between vehicles. The $ER$ denotes the effective observation range used to calibrate the sensor deployment, optimizing cost and performance.
@@ 3.1.2. Defuzzification: denoising function and fuzzy rules for TSP
where roundround signifies rounding up to the nearest integer. Based on $X_{defuzi}$, the number of waiting vehicles for each phase $WV_{p_{i}}$ is calculated, and fuzzy rules (detailed in Algorithm 1 lines 7-9) are used to determine the phase $p_{dynamic}$ corresponding to the largest number of waiting vehicles.
@@ 3.2.1. Fuzzification: Neural network fuzzy membership function (NNFM)
Lane Embedding Layer: First, we extract the number of vehicles of different segments (seg=ksliceseg=k*slice) of each traffic movement by recovered vehicle matrix $\hat{X}$. Where $k$ defines the length of the segment within each lane. To allow more capacity in feature extraction, each segment of the lane is embedded from one-dimensional into a higher dimensional latent space using a layer of multi-layer perception (MLP):
@@
(12) $t_{duration}=clip\left(h_{fuzzi}\ast refer\ duration+\epsilon,low,high\right)\\$
Where the $OU$ noise (Lillicrap et al., 2015) $\epsilon$ is used for enhancing phase duration exploration. The referrefer durationduration is calculated by maximum lane length $MaxLaneLength$ divided by maximum lane speed $MaxLaneSpeed$: $refer\ duration=\frac{MaxLaneLength}{MaxLaneSpeed}$.
We clip the phase duration to a maximum and minimum value to guarantee safety. To ensure pedestrian safety, set the pedestrian crossing time as the minimum duration of green lights lowlow.
@@
5:  for Each episode in epochsepochs do
@@ 5. Discussion
Scale
(Max_scale)
JiNan HangZhou
@@
Scale
(Max_scale)
JiNan HangZhou
@@ 2. How the FuzzyLight affects the throughput rate?
where $t_{duration}=\frac{dis}{v}$, disdis represents the distance of the last vehicle from intersections and $v$ denotes the average speed of the vehicle. We Then the throughput rate of a phase pair can be expressed as
@@
if $x(l_{i})=x(l_{j})$ then $Throughput=2$, therefore $1<=Throughput<=2$.